OpenStudy (anonymous):

Brittany's scores on her social studies tests are 82, 86, 95, 82, and 90. What is the mean of Brittany's scores?

OpenStudy (anonymous):

The mean is the average of the numbers, which you get by adding all of the numbers together and dividing based upon how many numbers there are. I think you can handle it from there with a calculator.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The mean is the average number. SUPER easy to do, just add up the numbers, then divide by how many numbers there are. SO, 82+86+95+82+90 = 435/5 = 87 :D

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

can i have a medal?

OpenStudy (anonymous):

im striving for 50 medals

OpenStudy (anonymous):

Uhh sure, I guess XD